OpenStudy (anonymous):

6 (4/5) - 3 (2/3)=? how to solve mixed fractions

OpenStudy (anonymous):

6(4/5)-3(2/3) ((6*5)+4))/5-((3*3)+2))/3 34/5-11/3 ((34*3)-(11*5))/15 [as lcm of 5 and 3 is 15] (102-55)/15 47/15

jhonyy9 (jhonyy9):

24/5 - 6/3 =24/5 - 2 = (24-10)/5 =14/5